TORSIONAL EXCITATION IN O-H STRETCH OVERTONE SPECTRA OF ETHYL HYDROPEROXIDE CONFORMERS

\underline{SHIZUKA HSIEH}, MA THIDA, MARGARET NYAMUMBO, and R. G. LINCK, Chemistry Department, Smith College, Northampton, MA 01063.

Photoacoustic spectra at 3-6 quanta of O-H stretch show features attributed to torsion about the O-O bond and to distinct contributions from two conformers. Laser-induced fluorescence detection of OH radicals demonstrates unimolecular dissociation from some vibrationally and torsionally excited states.